OpenStudy (anonymous):

Jon decides to take his seven year old son to the movies. The price for the child's ticket is 3.75 dollars less than the price for the adult's ticket. If you represent the price for the child's ticket using the variable "x", how would you write the algebraic expression for the adult's ticket price?

OpenStudy (anonymous):

The price for the child's ticket is 3.75 dollars less than the price for the adult's ticket, so if the price of the child's ticket is x and the adult ticket is y, then \[ x = y-3.75\] which gives us \[y = x+3.75\] <answer>